# Evaluation of cosmogenic Ge-68 background in a high purity germanium detector via a time series fitting method

### v1

@@ 2.4 Evaluation of the minimum detection activity
The minimum detection activity (MDA) of a $\gamma$ spectrometer relys on the background level, the measured time, the analyzed characteristic peak, and the detection efficiency. If other factors are keeped constant, the decay of Ge-68 will leads to a decrease of background and an improvement of the MDA.
Here we consider a typical ”pair measurement” to evaluate the improvement of MDA by the decrease of Ge-68 background. In a pair measurement, the sample and background are measured for the same time ($t_{m}$), and the MDA is written as[19]:
@@
where $i$ indicates the selected characteristic peak, $t_{U}$ is the underground operation time before the sample measurement, $I_{\gamma}$ is the yield of the characteristic $\gamma$ line, $\varepsilon$ is the detection efficiency, $b_{i}(t_{U},t_{m})$ is the background count in the analysis window ($E_{i}\pm 3\sigma_{E_{i}}$), $E_{i}$ is the energy of thecharacteristic $\gamma$ line $\sigma_{E_{i}}$ is the energy resolution. Background $b_{i}(t_{U},t_{m})$ is calculated using the 90 days background spectrum and the fitted Ge-68 activity ($A_{0,\mathrm{Ge-68}}$) via:
@@ 3.1 Detection efficiency of Ge-68 and radon daughter Bi-214
Total $10^{7}$ and $10^{8}$ particles are simulated for calculating the detection efficiencies of Ge-68 and airborne Bi-214, the simulated spectra are shown in Fig.3. For airborne Bi-214, its detection efficiencies in 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV energy ranges are 3.1, 0.73, and 7.6 cpd/(Bq/m3), respectively. For Ge-68, the efficiencies are 0.71, 1.20, and 132.3 cpd/(mBq/kg), respectively.
The count rate from 1 mBq/kg Ge-68 is equal to that from 17.4 Bq/m3 airborne Bi-214 in 1-3 MeV energy region. And the saturated activity of Ge-68 at sea level is around 1 mBq/kg[10]. This indicates that despite the spectrum of Ge-68 has no significant characteristic peak, it still can contribute a significant background in the total spectrum.
Figure 3: The simulated spectra of 1 mBq/kg Ge-68 and 1 Bq/m3 radon daughter Bi-214. The Ge-68 spectrum includes contribution from its decay daughter Ga-68 under equilibrium assumption. The three selected energy windows are labeled in gray.
@@ 3.2 Fitting result of Ge-68 activity
Figure 4: The best fit result of count rate in 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV. The best fit curves are labeled in black, the fitted contributions from Ge-68, Bi-214, and time-invariant background are labeled in red, green, and gray respectively. The residuals between the best fit and measured data are shown below the best fit curves.
The significance of Ge-68 signal is tested using method described in Sec.2.3. The calculated PDF of the test statistic $f(s)$ for null and alternative hypothesis are shown in Fig.5 alone with the observed value ($s_{obs}$). the $P$ value of the null hypothesis $P(H_{0})$=0.036%, which excludes the no Ge-68 hypothesis at 99.64% confidence level. The $P$ value of the alternative hypothesis $P(H_{1})$=33%, demonstrates the goodness of the fit and indicates a significant Ge-68 signal in the measured data.
@@
Figure 6: Contours for pairs of parameters in the $UltraNest$ MCMC sampling, and the projected likelihood distribution for each parameter. The colors indicates the relative value of the likelihood distribution, each vertical dash lines indicates the median position. For simplicity, the Bi-214 concentrations at different times are combined to an average value.
The fitted Ge-68 initial activity is 477.0$\pm$112.4 $\mu$Bq/kg, the uncertainty corresponding to the 68% confidence interval in the likelihood distribution drived from the $UltraNest$ MCMC sampling. Values and uncertainties of other parameters are listed in Tabel.1. Fig.6 demonstrates the contours for pairs of parameters in the MCMC sampling, which shows the correlation between each pairs of parameters. The largest correlation is between Ge-68 activity and time-invariant background in 1-3 MeV, which is also the main contributor of the uncertainty of Ge-68 activity.
@@ 3.3 Background induced by Ge-68 and its effects on MDA
Figure 7: Comparsion of measured spectrum (gray line) and simulated Ge-68 (Ga-68) spectrum (red line) with fitted activity. The fitted initial activity of Ge-68 is 477.0$\pm$112.4 $\mu$Bq/kg, the red shadow is the spectra corresponding to the fit uncertainty.
We select four characteristic $\gamma$ lines from different isotopes to evaluate the effects of Ge-68 background on their minimum detection activity (MDA). The select peaks are: 583.2 keV (Tl-208), 661.7 keV (Cs-137), 1460.8 keV (K-40), and 1764.5 keV (Bi-214). The 90 days background spectrum of the four characteristic $\gamma$ peaks and the contribution from Ge-68 (Ga-68) are shown in the left panel of Fig.8. For simplicity, our analysis assumes a constant background rate from the airborne radon daughter to better demonstrate the effects from the change of Ge-68 background. The improvement of MDA at different operation time in CJPL is calculated using eq.10, the results are shown in the right panel of Fig.8.
Figure 8: Left: the 90 days background spectrum in the four characteristic peak regions, contributions from Ge-68 (Ga-68) are labeled in red, the measurement of the background spectrum started about 21 days after the detector arrived CJPL. Right: the ratio of MDA between $t_{U}$ and $t_{U}=0$ for the four characteristic peaks.
After 5 years operation at CJPL, the Ge-68 activity decreases from 477.0 $\pm$112.4 $\mu$Bq/kg to 4.47$\pm$1.05 $\mu$Bq/kg and results in a 2$\sim$8MDA improvement for the selected four characteristic peaks. The smallest 2% improvement is for the Tl-208 583.2 keV peak as the background is dominated by Tl-208 in detector structure materials, Ge-68 only contributes 4.6% of the total background at $t_{U}$=0. For the Bi-214 1764.5 keV characteristic peak, Ge-68 contributes 21% of the total background at $t_{U}$=0 and 0.2% at $t_{U}$=5 years, the about 20% background reduction gives a 8% improvement in the MDA.
@@ 3.4 Variation of airborne Bi-214 concentration
Our method also provides the concentration variation of radon daughter Bi-214 in the detector chamber. As the detector chamber within the copper shielding is constantly purged by nitrogen gas, a comparsion between the variation of Bi-214 in detector chamber and radon in experiment hall indicates whether air has been mixed in the nitrogen gas or there is a leakage point in the shielding. The radon (Rn-222) concentration in the experiment hall has been measured by an AlphaGuard PQ2000 radon monitor. The AlphaGuard is set to the diffusion mode and a measurement period of one hour. Fig.9 demonstrates the comparsion of the Bi-214 and the Rn-222 concentration.
The variation of airborne Bi-214 is within 5 Bq/m3, its average value (1.42$\pm$0.17 Bq/m3) is about 40 times lower than the average Rn-222 concentration in experiment hall. And there is also no significant coincidence between the Bi-214 and Rn-222 concentration, indicating that the detector chamber has been well isolated from the the experiment hall.
Figure 9: Variation of the radon concentration in experiment hall and the airborne Bi-214 in detector chamber within the copper shielding. The gray point and black line are the radon concentration measured by an AlphaGuard radon monitor. The red line is the fitted airborne Bi-214 concentration and the red shadow is the corresponding fit uncertainty.
@@ 4 Conclusions
In this work, we develop a time series fitting method to calculate the activity of cosmogenic Ge-68 in a coaxial HPGe detector operated at China Jinping underground laboratory. Our method using the change of count rate in 1-3 MeV energy region and characteristic peaks of airborne radon daughter to decouple the Ge-68, radon daughter, and time-invariant background and fit the Ge-68 activity. The simulated detection efficiencies in different energy region are used to connect the count rate to Ge-68 activity and airborne Bi-214 concentration. Total 90 days measured data are used in the analysis and the Ge-68 initial activity is fitted to be 477.0$\pm$112.4 $\mu$Bq/kg, it contributes about 62% background in the 1-3 MeV energy region. Based on the measured background spectrum and fitted Ge-68 activity, we predict the minimum detection activity for four radioisotopes (Tl-208, Cs-137, K-40, and Bi-214) will improve by 2$\sim$8after 5 years underground operation.
Our method can be extended to other cosmogenic isotopes in germanium, for instance, Mn-54, Co-57, and Co-58. And the fitting result of Ge-68 activity could be used as an input in a spectrum fitting method to decouple background from different structure materials.

### v2

@@ 2.4 Evaluation of the minimum detection activity
The minimum detection activity (MDA) of a $\gamma$ spectrometer relies on the background level, the measured time, the analyzed characteristic peak, and the detection efficiency. If other factors are kept constant, the decay of Ge-68 will lead to a decrease in background and an improvement of the MDA.
Here we consider a typical ”paired measurement” to evaluate the improvement of MDA by the decrease of Ge-68 background. In a paired measurement, the sample and background are measured for the same time ($t_{m}$), and the MDA is written as[27]:
@@
where $i$ indicates the selected characteristic peak, $t_{U}$ is the underground operation time before the sample measurement, $I_{\gamma}$ is the yield of the characteristic $\gamma$ line, $\varepsilon$ is the detection efficiency, $b_{i}(t_{U},t_{m})$ is the background count in the analysis window ($E_{i}\pm 3\sigma_{E_{i}}$), $E_{i}$ is the energy of thecharacteristic $\gamma$ line $\sigma_{E_{i}}$ is the energy resolution. Background $b_{i}(t_{U},t_{m})$ is calculated using the 90-day background spectrum and the fitted Ge-68 activity ($A_{0,\mathrm{Ge-68}}$) via:
@@ 3.1 Detection efficiency of Ge-68 and radon daughter Bi-214
The detection efficiency is calculated via Geant4 simulation. To validate our simulation, the spectrum of the calibration experiment is compared with that obtained by the simulation. The calibration places three radioactive sources, Co-60, Cs-137, and Eu-152 at the top of the detector, The spectra of the three sources are simulated separately and then added together to compare with the measurement. As shown in in Fig.3. the simulated spectrum is in good agreement with the measured one.
Figure 3: Comparison between simulated and measured spectra in the calibration experiment. The calibration is performed with Co-60, Cs-137, and Eu-152, their simulated spectra are in green, blue, and orange, respectively. The red spectrum is the sum of the simulated spectra. The measured spectrum is labeled in black.
Total $10^{7}$ and $10^{8}$ particles are simulated for calculating the detection efficiencies of Ge-68 and airborne Bi-214, the simulated spectra are shown in Fig.4. For airborne Bi-214, its detection efficiencies in 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV energy ranges are 3.1, 0.73, and 7.6 cpd/(Bq/m3), respectively. For Ge-68, the efficiencies are 0.71, 1.20, and 132.3 cpd/(mBq/kg), respectively. The statistical uncertainties in the simulated detection efficiencies are less than 0.5% and therefore omitted in the analysis. The dead layer at the detector surface is set to 0.5 mm according to the manufactory, and its uncertainty is not considered in this work.
The count rate from 1 mBq/kg Ge-68 equals that from 17.4 Bq/m3 airborne Bi-214 in 1-3 MeV energy region. The saturated activity of Ge-68 at sea level is around 1 mBq/kg[10]. This indicates that despite the spectrum of Ge-68 has no significant characteristic peak, it still can contribute a significant background in the total spectrum.
Figure 4: The simulated spectra of 1 mBq/kg Ge-68 and 1 Bq/m3 radon daughter Bi-214. The Ge-68 spectrum includes the contribution from its decay daughter Ga-68 under the equilibrium assumption. The three selected energy windows are labeled in gray.
@@ 3.2 Fitting result of Ge-68 activity
Figure 5: The best-fit result of the count rate in 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV. The best-fit curves are labeled in black, and the fitted contributions from Ge-68, airborne Bi-214, and time-invariant background are labeled in red, green, and gray respectively. The residuals between the best-fit and measured data are shown below the best fit curves.
The significance of the Ge-68 signal is tested using the method described in Sec.2.3. The calculated PDF of the test statistic $f(s|H_{i})$ for null and alternative hypothesis are shown in Fig.6 alone with the observed value ($s_{obs}$). the $P$ value of the null hypothesis $P(H_{0})$=0.036%, which excludes the no Ge-68 hypothesis at 99.64% confidence level. The $P$ value of the alternative hypothesis $P(H_{1})$=33%, demonstrates the goodness of the fit and indicates a significant Ge-68 signal in the measured data.
@@
Figure 7: Contours for pairs of parameters in the $UltraNest$ MCMC sampling, and the projected likelihood distribution for each parameter. The colors indicate the relative value of the likelihood distribution, each vertical dash line indicates the median position. For simplicity, the Bi-214 concentrations at different times are combined to an average value.
The fitted Ge-68 initial activity is 477.0$\pm$112.4 $\mu$Bq/kg, the uncertainty corresponding to the 68% confidence interval in the likelihood distribution derived from the $UltraNest$ MCMC sampling. Values and uncertainties of other parameters are listed in Tabel.1. Fig.7 demonstrates the contours for pairs of parameters in the MCMC sampling, which shows the correlation between each pair of parameters. The largest correlation is between Ge-68 activity and time-invariant background in 1-3 MeV, which is also the main contributor to the uncertainty of Ge-68 activity.
Theoretical calculations give Ge-68 production rate in a range of 80-200 atom/kg/day at sea level for natural germanium detector [10, 28] corresponding to saturated activities of 925-2200 $\mu$Bq/kg. The uncertainly in the theoretical calculations mainly arise from the uncertainty of cosmic-ray neutron flux, which is affected by factories including altitude, latitude, and shielding. Our measured Ge-68 activity ($476.95\pm 112.38$ $\mu$Bq/kg) is significantly lower than the saturated activity at sea level. During the non-working hours in its fabrication, the detector was temporarily stored in a shallow underground location with an overburden exceeding 50 meters of water equivalent. The underground location provided shielding against the cosmic-rays and reduced the production rate of Ge-68 in the detector. However, due to the unknow exposure history before the detector left the manufactory, we cannot calculate the production rate of Ge-68 using the measured Ge-68 activity. For readers interested in the aclculation and evaluation of Ge-68 production rate, they can refer to[10, 28, 29, 30, 31] for more details.
@@ 3.3 Background induced by Ge-68 and its effects on MDA
Figure 8: Comparison of measured spectrum (gray line) and simulated Ge-68 (Ga-68) spectrum (red line) with fitted activity. The fitted initial activity of Ge-68 is 477.0$\pm$112.4 $\mu$Bq/kg, and the red shadow is the spectra corresponding to the fit uncertainty.
We select four characteristic $\gamma$ lines from different isotopes to evaluate the effects of Ge-68 background on their minimum detection activity (MDA). The select peaks are: 583.2 keV (Tl-208), 661.7 keV (Cs-137), 1460.8 keV (K-40), and 1764.5 keV (Bi-214). The 90-day background spectrum of the four characteristic $\gamma$ peaks and the contribution from Ge-68 (Ga-68) are shown in the left panel of Fig.9. For simplicity, our analysis assumes a constant background rate from the airborne radon daughter to better demonstrate the effects from the change of Ge-68 background. The improvement of MDA at different operation time in CJPL is calculated using Eq.10, the results are shown in the right panel of Fig.9.
Figure 9: Left: the 90-day background spectrum in the four characteristic peak regions, contributions from Ge-68 (Ga-68) are labeled in red, the measurement of the background spectrum started about 21 days after the detector arrived at CJPL. Right: the ratio of MDA between $t_{U}$ and $t_{U}=0$ for the four characteristic peaks.
After 5-year operation at CJPL, the Ge-68 activity decreases from 477.0 $\pm$112.4 $\mu$Bq/kg to 4.47$\pm$1.05 $\mu$Bq/kg and results in a 2$\sim$8MDA improvement for the selected four characteristic peaks. The smallest 2% improvement is for the Tl-208 583.2 keV peak as the background is dominated by Tl-208 in detector structure materials, Ge-68 only contributes 4.6% of the total background at $t_{U}$=0. For the Bi-214 1764.5 keV characteristic peak, Ge-68 contributes 21% of the total background at $t_{U}$=0 and 0.2% at $t_{U}$=5 years, the about 20% background reduction gives an 8% improvement in the MDA.
@@ 3.4 Variation of airborne Bi-214 concentration
Our method also provides the concentration variation of radon daughter Bi-214 in the detector chamber. As the detector chamber within the copper shielding is constantly purged by nitrogen gas, a comparison between the variation of Bi-214 in detector chamber and radon in experiment hall indicates whether air has been mixed in the nitrogen gas or there is a leakage point in the shielding. The radon (Rn-222) concentration in the experiment hall has been measured by an AlphaGuard PQ2000 radon monitor. The AlphaGuard is set to the diffusion mode and a measurement period of one hour. Fig.10 demonstrates the comparison of the Bi-214 and the Rn-222 concentration.
The variation of airborne Bi-214 is within 5 Bq/m3, and its average value (1.42$\pm$0.17 Bq/m3) is about 40 times lower than the average Rn-222 concentration in the experiment hall. There is also no significant coincidence between the Bi-214 and Rn-222 concentrations, indicating that the detector chamber has been well isolated from the experiment hall.
Figure 10: Variation of the radon concentration in the experiment hall and the airborne Bi-214 in the detector chamber within the copper shielding. The gray point and black line are the radon concentration measured by an AlphaGuard radon monitor. The red line is the fitted airborne Bi-214 concentration and the red shadow is the corresponding fit uncertainty.
3.5 Effects of time-invariant background on fitting sensitivity
The statistical fluctuations of the time-invariant backgrounds could jeopardize the time series fitting. Therefore, for other HPGe spectrometers with higher background levels, the sensitivity of our method could be degarded by the time-invariant backgrounds. To evaluate this effect, we tune our best fit model to generate count rate data with different time-invariant backgrounds. The count rate data are generated using Eq.2.2 and the best fit values in Table.1, except that the time-invariant backgrounds are multiply by a factor ($X$). Then the generated data are fitted by the time-series fitting method and the P-value of the null hypothesis ($P(H_{0})$) is calculated using Eq.6. Fig.11 shows the fit uncertainty of Ge-68 activities and $P(H_{0})$ as a function of the factor ($X$). As the time-invariant background increases, although the fitting parameters are still unbiased estimators, the uncertainty and P(H0) both increases, and when the factor $X$=4, corresponding to a $B_{1-3MeV}$ of 103.2 cpd, the significance of Ge-68 signal is less than 90%. It demonstrates that the control of the time-invariant background is a key to improve the sensitivity of our method.
Figure 11: The uncertainty of fitted Ge-68 activity (Left) and the P-value of the null hypothesis (Right) as a function of the factor ($X$). The time-invariant background is multiplied by $X$.
@@ 4 Conclusions
In this work, we develop a time series fitting method to calculate the activity of cosmogenic Ge-68 in a coaxial HPGe detector operated at China Jinping underground laboratory. Our method uses the change of count rate in 1-3 MeV energy region and characteristic peaks of airborne radon daughter to decouple the Ge-68, radon daughter, and time-invariant background and fit the Ge-68 activity. The simulated detection efficiencies in different energy regions are used to connect the count rate to Ge-68 activity and airborne Bi-214 concentration. A total 90-day measurement data are used in the analysis and the Ge-68 initial activity is fitted to be 477.0$\pm$112.4 $\mu$Bq/kg, it contributes about 62% background in the 1-3 MeV energy region. Based on the measured background spectrum and fitted Ge-68 activity, we predict the minimum detection activity for four radioisotopes (Tl-208, Cs-137, K-40, and Bi-214) will improve by 2$\sim$8after 5-year underground operation. For other low background experiments, for instance, experiment using HPGe to search for the Ge-76 neutrinoless double beta decay, Ge-68 is also an important background in the 2 MeV signal region[28]. Therefore, the sensitivity of such experiments could be benefit from the reduction of Ge-68 background. There are other measures to reduce the Ge-68 background in HPGe detectors, for instance, optimizing above-ground exposure time, additional shielding during transportation, and underground detector fabrication[17, 29, 30].
Our method can be extended to other cosmogenic isotopes in germanium, for instance, Mn-54, Co-57, and Co-58. As our method relies on observing a significant decay signal of the cosmogenic isotopes during the measurement time, it will be difficult to measure isotopes with a long half-life (for instance, H-3 and Co-60) using the time-series fitting method. The fitting result of Ge-68 activity could be used as an input to a spectrum fitting method to decouple background from different structure materials.